A lot of 12 Compresser tanks is checked to see whether there are any defective tanks. three thanks are checked for leaks. if 1 or more of the 3 is defective, the lot is rejected. Find the probability that the lot will be rejected if three are actually 3 defective thanks in the lot.


1
Expert's answer
2021-12-29T14:34:13-0500

X has the hypergeometric distribution:


"\ud835\udc4e = 3 (\ud835\udc51\ud835\udc52\ud835\udc53\ud835\udc52\ud835\udc50\ud835\udc61\ud835\udc56\ud835\udc63\ud835\udc52), \ud835\udc4f = 9 (\ud835\udc52\ud835\udc53\ud835\udc53\ud835\udc52\ud835\udc50\ud835\udc61\ud835\udc56\ud835\udc63\ud835\udc52)"

"𝑋 = 0 (𝑛𝑜𝑛𝑒 \ 𝑑𝑒𝑓𝑒𝑐𝑡𝑖𝑣𝑒)"

"n=3, \ud835\udc5b \u2212 \ud835\udc4b = 3"

"P(X)=\dfrac{\dbinom{a}{X}\dbinom{b}{n-X}}{\dbinom{a+b}{n}}"

"=\dfrac{\dbinom{3}{0}\dbinom{9}{3-0}}{\dbinom{3+9}{3}}"

"=\dfrac{1(84)}{220}=\dfrac{21}{55}"

The probability that the lot will be rejected is


"1-\dfrac{21}{55}=\dfrac{34}{55}\approx0.6182"